Vertical Co. has a debt to equity ratio of 0.50.  The company is considering a new plant that will cost $150 million to build. When the company issues new equity, it incurs a flotation cost of 7.5%. The flotation cost on new debt is 2%.

a) What are the weighted average flotation costs? Enter percentages as decimals and round to 4 decimals

b) What is the cost of the plant, including flotation costs? Enter percentages as decimals and round to 4 decimals
Answer #1

a. Weighted average flotation cost =Weight of Equity*Cost of flotation of Equity +Weight of Debt*Cost of flotation of debt
=1/(1+0.5)*7.5%+0.5/(1+0.5)*2% =0.056667 or 0.0567

b. Cost of Plant =New Plant cost +New Plant Cost*0.0567 =150+150*0.056667 =150+8.50 =158.50